# AR Surgical Navigation With Surface Tracing: Comparing In-Situ Visualization with Tool-Tracking Guidance for Neurosurgical Applications

### v1

@@ 4 Discussion
4.1 Influence of Real‑Time Tool Tracking on Accuracy
Real‑time tool tracking yielded pronounced accuracy gains across nearly every metric. Most notably, median target–tip error fell from 17.5 mm to 4.3 mm, and depth error from 14.5 mm to 2.2 mm, while lateral deviation and angular error were each roughly halved. Catheter entry offset improved modestly but not significantly, and signed depth error exhibited a non‑significant trend in favour of tool-tracking. These results corroborate our hypothesis that continuous six‑degree‑of‑freedom feedback mitigates both angular drift and excessive insertion, leading to more precise ventricular targeting. Much of the improvement stems from reduced angular error. Small deviations in trajectory quickly compound into large target offsets over the catheter’s length. Tool tracking helps users correct these in real time, unlike static overlays.
@@ 5 Conclusion
We presented an augmented‑reality neurosurgical navigation system that combines landmark‑ and surface‑based registration with real‑time tool tracking on a head‑mounted display. In a controlled phantom study, the approach cut median target–tip error by more than fourfold (4.3 mm vs 17.5 mm) and reduced overshoot from 14.5 mm to 2.2 mm, while also halving angular and lateral deviations—all without materially lengthening the procedural workflow. Participants reported higher confidence and lower cognitive load under tool tracking compared with static in‑situ visualisation. These findings underscore the promise of compact, easily deployed AR guidance for ventricular catheter placement and motivate continued evaluation in live clinical environments.

### v2

@@ 4 Discussion
4.1 Influence of Real-Time Tool Tracking on Accuracy
Real-time tool tracking yielded pronounced accuracy gains across nearly every metric. Most notably, median target–tip error fell from 17.5 mm to 4.3 mm, and depth error from 14.5 mm to 2.2 mm, while lateral deviation and angular error were each roughly halved. Catheter entry offset improved modestly but not significantly, and signed depth error exhibited a non-significant trend in favour of tool-tracking. These results corroborate our hypothesis that continuous six-degree-of-freedom feedback mitigates both angular drift and excessive insertion, leading to more precise ventricular targeting. Much of the improvement stems from reduced angular error. Small deviations in trajectory quickly compound into large target offsets over the catheter’s length. Tool tracking helps users correct these in real time, unlike static overlays.
@@ 5 Conclusion
We presented an augmented-reality neurosurgical navigation system that combines landmark- and surface-based registration with real-time tool tracking on a head-mounted display. In a controlled phantom study, the approach cut median target–tip error by more than fourfold (4.3 mm vs 17.5 mm) and reduced overshoot from 14.5 mm to 2.2 mm, while also halving angular and lateral deviations—all without materially lengthening the procedural workflow. Participants reported higher confidence and lower cognitive load under tool tracking compared with static in-situ visualisation. These findings underscore the promise of compact, easily deployed AR guidance for ventricular catheter placement and motivate continued evaluation in live clinical environments.